# Gravitational Atoms from Topological Stars

### v1

@@ II Massive bound states of Topological Stars
parametrized by a local real function $\alpha(x)$. Here, $\mu$ is the mass of the field and $e$ its Maxwell coupling constant, or fundamental charge. The TS-black string solution Bah and Heidmann (2021a) is a 2-parameter family with vanishing scalar, $\Phi=0$, given by
@@
where $r_{s,b}$ are free parameters and the magnetic charge $P$ is fixed to:
@@
so $K$ labels one-parameter families of solutions. The family with $K=1$ is smooth at $r=r_{b}$ and the global space topology is $R^{2}\times S^{2}$, while $K>1$ describes an orbifold geometry $R^{2}/\mathbb{Z}_{K}\times S^{2}$. Fluctuations about those backgrounds include metric $\delta g_{\mu\nu}$, electromagnetic $\delta A_{\mu}$ and scalar perturbations $\delta\Phi$. However, due to the vanishing scalar background, first-order scalar fluctuations decouple from the gravitational and electromagnetic ones, and satisfy (henceforth we write $\delta\Phi\equiv\Phi$ to simplify notation),
@@
Our aim in this section is to establish some general properties of solutions to (10). This is similar to the case of neutral fields, previously considered Heidmann et al. (2023a); Bianchi et al. (2023), except for the angular analysis, which presents substantial differences. First, to ensure that solutions to (10) belong to a smooth $U(1)$-bundle, one needs to impose Dirac’s quantization condition Dirac (1931),
@@
Separable solutions to (10) can be constructed by employing the Wu-Yang monopole harmonics $Y_{N\ell m}(\theta,\phi)=\mathcal{P}_{N,\ell,m}(\theta)e^{im\phi}$ Wu and Yang (1976) (we review them in Appendix A),
@@
where $M$ denotes the spacetime manifold. This shows that only the zero-KK modes are smooth functions, a fact that can be interpreted as the scalar field probing a delta-source at the star’s center. However, as we will see below, the conserved physical currents associated to $\Phi$ (e.g. the energy) are first-derivative functionals of $\Phi$, so the fact that in all cases the differentiability class of $\Phi$ is $C^{p\geq 1}(M)$ guarantees that (17) indeed corresponds to physically-regular solutions.
At infinity, solutions can present two linearly-independent behaviors, $\psi(r)\sim e^{\pm i\sqrt{\omega^{2}-k^{2}-\mu^{2}}\ r}$, and requiring that only one of them is present,
@@
defines a characteristic-value problem for the frequencies, $\omega$, which in general will allow only a discrete set of solutions. Assuming $\omega$ satisfies such problem, the solutions fall into two classes:
QNMs: $\quad\text{Im}\left[\sqrt{\omega^{2}-k^{2}-\mu^{2}}\right]<0$ (20)
Bound states: $\quad\text{Im}\left[\sqrt{\omega^{2}-k^{2}-\mu^{2}}\right]>0\,$
@@
In general, in BH spacetimes bound states decay fast in time, being absorbed by the hole. However, in the case of rotating BHs bound states can grow in time due to superradiance, a process known as superradiant instability Brito et al. (2015a). At the threshold separating time-growing and time-decaying solutions lie the stationary states, which are bound to the hole and oscillate at a specific frequency (the superradiant frequency), but do not decay in time. This yields the notion of a gravitational atom, discussed in the introduction.
In the case of TSs, bound states are not only unable to leak out at infinity, but also lack a dissipation mechanism at the star surface, due to the event horizon being replaced by a perfectly reflecting boundary. Indeed, in the next section we show that bound states in TS spacetimes are strictly normal modes, that is, they oscillate in time, but do not decay. These give rise to a genuine realization of a gravitational atom, superseding systems involving BHs in the sense that they are stable configurations–the only dissipation mechanism is through the emission of gravitational and electromagnetic waves. In addition, these states turn out to exhibit a rich structure that we discuss at length in Section III.
Before that, here we establish some generic aspects. Namely, we first prove that any bound state of a TS must be strictly a normal mode. Next, we show that in the high-frequency regime, where the field’s Compton wave length $\lambda=1/\mu$ is short compared to the TS’s size, $\mu r_{b}\gg 1$, normal modes are described by congruences of timelike geodesics. This allows us to identify two broad classes of modes, dubbed cloud modes and inner modes, with qualitatively distinct properties. Finally, we show that in the opposite regime, where $\mu r_{b}\ll 1$, the spectrum becomes hydrogenic. This is similar to the states of the BH gravitational atom Detweiler (1980b); Baumann et al. (2019a) but, in our case, the clouds are stable and non-decaying due to the absence of an event horizon.
@@ II.1 Proving that bound modes are normal modes
Heuristically, bound modes of topological stars should persist over time due to energy conservation, as being exponentially suppressed at infinity and in the absence of horizons there is no available channel for their dissipation. An elegant way to prove this is to employ the conservation laws associated to fields propagating on the background (II), as we do next. Alternatively, one could simply refer to the hermiticity properties of the ODE problem satisfied by $R(r)$ in (12), in the case of bound states. We prefer the former approach, since we find it more geometric and physically insightful and does not rely on unphysical statements such as the separability properties of the spatial coordinates.
@@
Then, if $\Phi$ satisfies its equation of motion, the following local functional [^3] Here, the notation $\boldsymbol{\epsilon}$ is the volume form and the notation $\boldsymbol{\epsilon}_{\mu}T^{\mu}$ stands for $\boldsymbol{\epsilon}_{\mu\mu_{1}...\mu_{d-1}}T^{\mu}$, in contrast with our conventions for the hodge dual, $\star T_{\mu_{1}...\mu_{d-1}}=\boldsymbol{\epsilon}_{\mu_{1}...\mu_{d-1}\nu}T^{\nu}$. To avoid cluttering notation, we do not write the indices of differential forms, and just use bold symbols to indicate that omission, following the notational conventions in Wald (1993). is conserved:
@@
The conservation of (22) on-shell follows by construction from the gauge symmetries of the theory (see Elgood et al. (2020); Ortin and Pereñiguez (2022); Dyson and Pereñiguez (2023)), but can also be verified explicitly using the Noether identities given in Appendix B. For the TS (II), the Killing vectors $\partial_{t},\partial_{\phi}$ and $\partial_{y}$ have associated the momentum maps,
@@
and the associated conserved functionals can be interpreted as changes in ADM mass, angular- and KK momentum due to the presence of the scalar field Ortin and Pereñiguez (2022); Dyson and Pereñiguez (2023); Ballesteros et al. (2023); Gomez-Fayren et al. (2023). It will suffice to consider the energy functional associated to $\partial_{t}$. We shall evaluate it on a mode of the form (12), assuming that the frequency possesses both real and imaginary parts,
@@
Then, integrating it on a constant time slice of (II), denoted $\Sigma_{t}$ (see Figure 3), gives
@@
where we assumed the orbifold condition (9) with $K=1$ for simplicity, the smooth TS, but the argument can be extended to the general background with $K\geq 1$.
@@
The fact that the modes are at least $C^{1}(M)$ in the vicinity of $r=r_{B}$, together with the bound-state condition in (20), guarantee that the integral is finite. Then conservation of $E[\Phi]$ implies that it cannot depend on the time slice chosen, [^4] One subtlety is that, for Stokes’ theorem to apply in this case, modes should be at least $C^{2}(M)$ functions, which is not true for modes carrying precisely one unit of KK momentum, see (18). In that case, one can modify the integration domain adding a timelike boundary at $r=r_{B}+\delta$, apply Stokes theorem, and show easily that the additional integral vanishes as $\delta\to 0$. The same strategy can be used to generalize our argument to the TS with non-trivial orbifolds $K>1$. so it must be that
@@
We can thus conclude that bound modes are strictly normal modes. The numerical and analytic results of the following sections confirm this fact, as expected. A similar strategy can be used in the case of BHs, where the integrals at two time slices would differ by an integral across the event horizon (see Pereñiguez et al. (2024) for the related case of a magnetic BH). From the latter, it follows immediately that a bound state that is not dissipating through the horizon necessarily oscillates at the superradiant frequency and does not decay in time.
II.2 The geometric particle limit $\mu r_{b}\gg 1$: inner v.s. cloud modes
A great deal of intuition about the structure of normal modes can be gained by studying a geometric particle limit. Similarly to the usual geometric optics limit, this focuses on fluctuations whose characteristic wavelength is much shorter than the background’s lengthscale. For simplicity, we will restrict ourselves to neutral modes, with no KK momentum. The idea consists in making a large-phase ansatz, assuming that the field’s Compton wavelength $\lambda=1/\mu$ is much shorter than the star’s size, $\lambda/r_{b}\ll 1$. Writing
@@
and inserting this into the wave equation and expanding in $\lambda/r_{b}$ one finds, to leading order,
@@
so in the small-$\lambda$ limit the solutions to the wave equation are governed by congruences of timelike geodesics. To solve equation (29) (restricting again to the case of zero KK momentum), we note that the principal function $S$ admits separable solutions of the form
@@
So far, $\mathcal{E}$, $\mathcal{M}$ and $\mathcal{Q}$ are arbitrary (except for the condition that $r_{b}\mathcal{M}/\lambda\in Z$, to guarantee that the solution is single-valued). For reasons that will be clear soon, we shall fix them by requiring that $\mathcal{R}(r)$ and $\Theta(\theta)$ vanish and have maxima at some radius $r=r_{0}$ and the equator $\theta=\pi/2$, respectively. That sets $\mathcal{Q}=0$, and fixes $\mathcal{E}$ and $r_{0}$ in terms of $\mathcal{M}$ as
@@
With these choices, the solutions to (32) for both $S_{\theta}(\theta)$ and $S_{R}(r)$ are purely imaginary, so the wave (34) is exponentially localized at the equator and at $r=r_{0}(\mathcal{M})$. At large values of $\mathcal{M}$, the frequency of the mode and the radial position at which it localizes are
@@
where the ellipsis denotes subleading terms in the large-$\mathcal{M}$ expansion. The modes localize closer to the star as $\mathcal{M}$ decreases. However, below a threshold value $\mathcal{M}_{min}=\sqrt{3}r_{s}/r_{b}$ the behavior changes qualitatively. For $\mathcal{M}=\mathcal{M}_{min}$, one has
$\omega=\frac{1}{\lambda}\frac{2\sqrt{2}}{3}=\frac{1}{\lambda}\mathcal{E}_{ISCO}\,,\quad r_{0}=3r_{s}\equiv r_{ISCO}\,$, (38)
where we used that the ISCO of TSs lies at $3r_{s}$. However, the mode now is not localized at $r=r_{ISCO}$, since that becomes an inflection point of $\mathcal{R}(r)$ rather than a maximum. We have $\mathcal{R}(r)>0$ for $r<r_{ISCO}$ and $\mathcal{R}(r)<0$ for $r>r_{ISCO}$, so the corresponding solution (34) oscillates within the region $r<r_{ISCO}$, but is exponentially suppressed beyond it.
This analysis predicts the existence of two broad classes of modes in the regime $\mu r_{b}\gg 1$. For large angular momentum $m\gg\mu r_{b}$ (so $\ell\gg\mu r_{b}$), the solutions to the massive wave equation are described by (34), with $\mathcal{E}$ given by (37) and the modes localizing far away from the star, at the radius predicted by (37). We shall refer to these as cloud modes. On the other hand, for lower angular momentum $\ell\ll\,r_{b}\mu$ modes can probe the star’s center, so we refer to them as inner modes. The boundary between cloud and inner modes occurs at $\ell\sim\sqrt{3}\mu r_{s}$, where the modes oscillate at approximately the ISCO energy, as predicted in (38). We expect any horizonless compact object to admit a similar mode classification in the small-$\lambda$ regime. However, for finite values of $\lambda$, the inner modes will exhibit properties that depend on the object’s structure. In the next section, we confirm our predictions at small $\lambda$, and establish a more accurate classification of the TS inner modes for arbitrary values of the parameters.

### v2

@@ II Massive bound states of Topological Stars
parametrized by a local real function $\alpha(x)$. Here, $\mu$ is the field mass and $e$ its Maxwell coupling constant, or fundamental charge. The TS-black string solution Bah and Heidmann (2021a) is a 2-parameter family with vanishing scalar, $\Phi=0$, given by
@@
where $r_{s,b}$ are free parameters and the magnetic charge $P$ is fixed to
@@
so $K$ labels one-parameter families of solutions. The family with $K=1$ is smooth at $r=r_{b}$ and the global space topology is $R^{2}\times S^{2}$, while $K>1$ describes an orbifold geometry $R^{2}/\mathbb{Z}_{K}\times S^{2}$. Fluctuations about those backgrounds include metric ($\delta g_{\mu\nu}$), electromagnetic ($\delta A_{\mu}$) and scalar ($\delta\Phi$) perturbations. However, due to the vanishing scalar background, first-order scalar fluctuations decouple from the gravitational and electromagnetic ones, and satisfy (henceforth we write $\delta\Phi\equiv\Phi$ to simplify notation),
@@
Our aim in this section is to establish some general properties of solutions of Eq. (10). This is similar to the case of neutral fields, previously considered Heidmann et al. (2023a); Bianchi et al. (2023), except for the angular analysis, which presents substantial differences. First, to ensure that solutions of Eq. (10) belong to a smooth $U(1)$-bundle, one needs to impose Dirac’s quantization condition Dirac (1931),
@@
Separable solutions of Eq. (10) can be constructed by employing the Wu-Yang monopole harmonics $Y_{N\ell m}(\theta,\phi)=\mathcal{P}_{N,\ell,m}(\theta)e^{im\phi}$ Wu and Yang (1976) (we review them in Appendix A),
@@
where $M$ denotes the spacetime manifold. This shows that only the zero-KK modes are smooth functions, a fact that can be interpreted as the scalar field probing a delta-source at the star’s center. However, as we will see below, the conserved physical currents associated to $\Phi$ (e.g., the energy) are first-derivative functionals of $\Phi$, so the fact that in all cases the differentiability class of $\Phi$ is $C^{p\geq 1}(M)$ guarantees that Eq. (17) indeed corresponds to physically regular solutions.
At infinity, solutions can present two linearly independent behaviors, $\psi(r)\sim e^{\pm i\sqrt{\omega^{2}-k^{2}-\mu^{2}}\ r}$, and requiring that only one of them is present,
@@
defines a characteristic value problem for the frequencies, $\omega$, which in general will allow only a discrete set of solutions. Assuming $\omega$ satisfies such problem, the solutions fall into two classes:
QNMs: $\quad\text{Im}\left[\sqrt{\omega^{2}-k^{2}-\mu^{2}}\right]<0\,$, (20)
Bound states: $\quad\text{Im}\left[\sqrt{\omega^{2}-k^{2}-\mu^{2}}\right]>0\,.$
@@
In general, in BH spacetimes bound states decay fast in time, being absorbed by the hole. However, in the case of rotating BHs bound states can grow in time due to superradiant instabilities Brito et al. (2015a). At the threshold separating time-growing and time-decaying solutions lie the stationary states, which are bound to the hole and oscillate at a specific frequency (the superradiant frequency), but do not decay in time Hod (2012); Herdeiro and Radu (2014). This yields the notion of a gravitational atom, discussed in the introduction.
In the case of TSs, bound states are not only unable to leak out at infinity, but also lack a dissipation mechanism at the star surface, due to the event horizon being replaced by a perfectly reflecting boundary. Indeed, in the next section we show that bound states in TS spacetimes are strictly normal modes, that is, they oscillate in time, but do not decay. These give rise to a genuine realization of a gravitational atom, superseding systems involving BHs in the sense that they are stable configurations – the only dissipation mechanism is through the emission of gravitational and electromagnetic waves. In addition, these states turn out to exhibit a rich structure, that we discuss at length in Section III.
Before that, here we establish some generic facts. Namely, we first prove that any bound state of a TS must be strictly a normal mode. Next, we show that in the high-frequency regime, where the field’s Compton wavelength $\lambda=1/\mu$ is short compared to the TS’s size, $\mu r_{b}\gg 1$, normal modes are described by congruences of timelike geodesics. This allows us to identify two broad classes of modes, dubbed cloud modes and inner modes, with qualitatively distinct properties. Finally, we show that in the opposite regime, where $\mu r_{b}\ll 1$, the spectrum becomes hydrogenic. This is similar to the states of the BH gravitational atom Detweiler (1980b); Baumann et al. (2019a) but, in our case, the clouds are stable and non-decaying due to the absence of an event horizon.
@@ II.1 Proving that bound modes are normal modes
Heuristically, bound modes of TSs should persist over time due to energy conservation, as being exponentially suppressed at infinity and in the absence of horizons there is no available channel for their dissipation. An elegant way to prove this is to employ the conservation laws associated to fields propagating on the background (II), as we do next. Alternatively, one could simply refer to the hermiticity properties of the ODE problem satisfied by $R(r)$ in Eq. (12), in the case of bound states. We prefer the former approach, since we find it more geometric and physically insightful and it does not rely on unphysical statements, such as the separability properties of the spatial coordinates.
@@
Then, if $\Phi$ satisfies its equation of motion, the following local functional [^2] Here, $\boldsymbol{\epsilon}$ is the volume form and the notation $\boldsymbol{\epsilon}_{\mu}T^{\mu}$ stands for $\boldsymbol{\epsilon}_{\mu\mu_{1}...\mu_{d-1}}T^{\mu}$, in contrast with our conventions for the hodge dual, $\star T_{\mu_{1}...\mu_{d-1}}=\boldsymbol{\epsilon}_{\mu_{1}...\mu_{d-1}\nu}T^{\nu}$. To avoid cluttering notation, we do not write the indices of differential forms, and just use bold symbols to indicate that omission, following the notational conventions in Wald (1993). is conserved:
@@
The conservation equation (22) on-shell follows by construction from the gauge symmetries of the theory (see Elgood et al. (2020); Ortin and Pereñiguez (2022); Dyson and Pereñiguez (2023)), but can also be verified explicitly using the Noether identities given in Appendix B. For the TS (II), the Killing vectors $\partial_{t},\partial_{\phi}$ and $\partial_{y}$ have associated the momentum maps,
@@
and the associated conserved functionals can be interpreted as changes in ADM mass, angular- and KK-momentum due to the presence of the scalar field Ortin and Pereñiguez (2022); Dyson and Pereñiguez (2023); Ballesteros et al. (2023); Gomez-Fayren et al. (2023). It will suffice to consider the energy functional associated to $\partial_{t}$. We shall evaluate it on a mode of the form (12), assuming that the frequency possesses both real and imaginary parts,
@@
Then, integrating it on a constant time slice of (II), denoted $\Sigma_{t}$ (see Fig. 3), gives
@@
where we assumed the orbifold condition (9) with $K=1$ for simplicity, i.e., the smooth TS, but the argument can be extended to the general background with $K\geq 1$.
@@
The fact that the modes are at least $C^{1}(M)$ in the vicinity of $r=r_{B}$, together with the bound-state condition in Eq. (20), guarantee that the integral is finite. Then conservation of $E[\Phi]$ implies that it cannot depend on the time slice chosen, [^3] One subtlety is that, for Stokes’ theorem to apply in this case, modes should be at least $C^{2}(M)$ functions, which is not true for modes carrying precisely one unit of KK momentum, see Eq. (18). In that case, one can modify the integration domain adding a timelike boundary at $r=r_{B}+\delta$, apply Stokes theorem, and show easily that the additional integral vanishes as $\delta\to 0$. The same strategy can be used to generalize our argument to the TS with non-trivial orbifolds $K>1$. so it must be that
@@
We can thus conclude that bound modes are strictly normal modes. The numerical and analytical results of the following sections confirm this fact, as expected. A similar strategy can be used in the case of BHs, where the integrals at two time slices would differ by an integral across the event horizon (see Pereñiguez et al. (2024) for the related case of a magnetic BH). From the latter, it follows immediately that a bound state that is not dissipating through the horizon necessarily oscillates at the superradiant frequency and does not decay in time.
II.2 The geometric particle limit $\mu r_{b}\gg 1$:
inner vs. cloud modes
A great deal of intuition about the structure of normal modes can be gained by studying a geometric particle limit. Similarly to the usual geometric optics limit, this focuses on fluctuations whose characteristic wavelength is much shorter than the background’s lengthscale. For simplicity, we will restrict ourselves to neutral modes, with no KK momentum. The idea consists in making a large-phase ansatz, assuming that the field’s Compton wavelength $\lambda=1/\mu$ is much shorter than the star’s size, i.e., $\lambda/r_{b}\ll 1$. Writing
@@ II.2 The geometric particle limit $\mu r_{b}\gg 1$:
inserting this into the wave equation and expanding in $\lambda/r_{b}$ one finds, to leading order,
@@
so in the small-$\lambda$ limit the solutions to the wave equation are governed by congruences of timelike geodesics. To solve Eq. (29) (restricting again to the case of zero KK momentum), we note that the principal function $S$ admits separable solutions of the form
@@
So far, $\mathcal{E}$, $\mathcal{M}$ and $\mathcal{Q}$ are arbitrary (except for the condition that $r_{b}\mathcal{M}/\lambda\in Z$, to guarantee that the solution is single-valued). For reasons that will be clear soon, we shall fix them by requiring that $\mathcal{R}(r)$ and $\Theta(\theta)$ vanish and have maxima at some radius $r=r_{0}$ and at the equator $\theta=\pi/2$, respectively. That sets $\mathcal{Q}=0$, and fixes $\mathcal{E}$ and $r_{0}$ in terms of $\mathcal{M}$ as
@@
With these choices, the solutions of Eq. (32) for both $S_{\theta}(\theta)$ and $S_{R}(r)$ are purely imaginary, so the wave (34) is exponentially localized at the equator and at $r=r_{0}(\mathcal{M})$. At large values of $\mathcal{M}$, the frequency of the mode and the radial position at which it is localized are
@@
where the ellipsis denotes subleading terms in the large-$\mathcal{M}$ expansion. The modes are localized closer to the star as $\mathcal{M}$ decreases. However, below a threshold value $\mathcal{M}_{\rm min}=\sqrt{3}r_{s}/r_{b}$ the behavior changes qualitatively. For $\mathcal{M}=\mathcal{M}_{\rm min}$, one has
$\omega=\frac{1}{\lambda}\frac{2\sqrt{2}}{3}=\frac{1}{\lambda}\mathcal{E}_{\rm ISCO}\,,\quad r_{0}=3r_{s}\equiv r_{\rm ISCO}\,$, (38)
where we used that the ISCO of TSs lies at $3r_{s}$. However, the mode now is not localized at $r=r_{\rm ISCO}$, since that becomes an inflection point of $\mathcal{R}(r)$ rather than a maximum. We have $\mathcal{R}(r)>0$ for $r<r_{\rm ISCO}$ and $\mathcal{R}(r)<0$ for $r>r_{\rm ISCO}$, so the corresponding solution (34) oscillates within the region $r<r_{\rm ISCO}$, but is exponentially suppressed beyond it.
This analysis predicts the existence of two broad classes of modes in the regime $\mu r_{b}\gg 1$. For large angular momentum $m\gg\mu r_{b}$ (so $\ell\gg\mu r_{b}$), the solutions of the massive wave equation are described by Eq. (34), with $\mathcal{E}$ given by Eq. (37) and the modes localized far away from the star, at the radius predicted by Eq. (37). We shall refer to these as cloud modes. On the other hand, for lower angular momentum $\ell\ll\,r_{b}\mu$ modes can probe the star’s center, so we refer to them as inner modes. The boundary between cloud and inner modes occurs at $\ell\sim\sqrt{3}\mu r_{s}$, where the modes oscillate at approximately the ISCO energy, as predicted in Eq. (38). We expect any horizonless compact object to admit a similar mode classification in the small-$\lambda$ regime. However, for finite values of $\lambda$, the inner modes will exhibit properties that depend on the object’s structure. In the next section, we confirm our predictions at small $\lambda$, and establish a more accurate classification of the TS inner modes for arbitrary values of the parameters.